OpenStudy (anonymous):

Multiply and simplify by factoring. Assume that all expressions under radicals represent nonneggative numbers. (see comments for work)

OpenStudy (anonymous):

\[\sqrt{18a^13b} \times \sqrt{27a^14b^12}\]

OpenStudy (anonymous):

the last part is supose to be 12 so its 4b^12